# Effects of rotation and anisotropy on the properties of type-II holographic superconductors

### v1

@@ 5 Turning on a uniform external magnetic field
with $\zeta=-\dfrac{a}{\Xi}\,\rho$. Regular boundary conditions—both asymptotically and at the horizon—constrain the parameters in the following manner: the critical exponent must lie within the interval $31\leq z<$ 3, while the constants are fixed as $\rho=r_{h}^{3-z}\mu$ and $\zeta=r_{h}^{3-z}\nu$.
@@
[…]
@@ 5.1 The external magnetic field
Upon a brief inspection of Eq. (48), it becomes evident that the external magnetic field plays a significant role in the dynamics of the radial mode. As we proceed to delve into this issue, the consistency of the solution unveils a relationship between the upper critical magnetic field and the black hole parameters.
The standard holographic prescription provides the renormalizable boundary form of the scalar field as given below
@@
where the scaling dimensions are $\Delta_{\pm}=\dfrac{z+3\pm\sqrt{(z+3)^{2}+4\ell^{2}m^{2}}}{2}$. Without loss of generality, we may impose the boundary condition $J_{-}=0$, and $J_{+}=J$ with scaling dimension $\Delta_{+}=\Delta$.
Similarly, we propose a near-horizon Taylor expansion of $R(u)$ around $u=1$, namely
@@
Evaluation of Eq. (48) permits to algebraically solve order by order the coefficients of the series in terms of $R(1)$
@@
Once again, the matching method described in Zhao:2013pva becomes relevant. Ensuring continuity and regularity requires a smooth matching between (49) and (50), and their derivatives, within a specific region of the bulk domain, say $u=u_{m}$. This matching can be achieved at the expense of the following algebraic constraints
@@
×$\times\Biggl{\{}\ell^{2}m^{2}\Big{[}\ell^{2}m^{2}+2(z+3)\Big{]}+\dfrac{\ell^{4}B}{r_{h}^{2}}\Big{[}2\ell^{2}m^{2}+\dfrac{\ell^{4}B}{r_{h}^{2}}\Big{]}-\quantity(\dfrac{\ell}{\Xi})^{2}\quantity(\dfrac{\ell}{r_{h}})^{2z}[{A_{t}^{(0)}}^{\prime}(1)]^{2}\Biggr{\}}$
×$\times\,R(1)(u_{m}-1)^{2},$ (52a)
@@
A most relevant implication arises from the evaluation of (52) at the upper critical magnetic field, $B\approx B_{c2}$, consistent with the perturbation parameter $\epsilon\ll 1$, and the subsequent insertion of the expression for ${A_{t}^{(0)}}^{\prime}(1)=(z-3)\dfrac{\rho}{r_{h}^{3-z}}$ obtained from (44). We are left with a consistency condition which can be read as a formula for the upper critical magnetic field in terms of the critical temperature, i.e.,
@@
A quick check-up reveals that $B_{c2}$ goes as predicted by the Ginzburg-Landau theory near the critical temperature, with $B_{c2}\propto(1-T/T_{c})$. This behavior is depicted in Figures 5(a) and 5(b).
Figure 5: Figure (a) shows the behavior of the upper critical magnetic field by setting $\ell^{2}m^{2}=-3$, $u_{m}=0.8$, and $z=2$ while taking diverse values of the rotation parameter. On the other hand, Figure (b) is obtained by setting $\ell^{2}m^{2}=-3$, $u_{m}=0.9$ and $a=1$ while considering different values of the critical exponent $z$.
@@ 5.2 The vortex configuration
As previously anticipated, our focus lies in developing a model for the vortex lattice deformation in measurable superconductors. At this stage, we can conclude the theory we have formulated and proceed to characterize the lattice properties, similar to the investigations carried out in previous works Maeda:2009vf ; Guo:2014wca ; Herrera-Mendoza:2022whz .
@@
The lattice structure is fully attained up to this point. It is straightforward to verify that the modulus squared |$\absolutevalue{\gamma_{L}}^{2}$ is already periodic while fundamental regions conforming the lattice are generated by the vectors (61). Consequently, any point on the lattice can be expressed as a linear combination of $\vec{b}_{1}$ and $\vec{b}_{2}$. In agreement with Ginzburg-Landau theory, here the magnetic flux is quantized as $B\times 2\pi r_{0}^{2}=2\pi$.
The zeros of |$\absolutevalue{\gamma_{L}}$ | indicate the locations of vortex cores, where the order parameter vanishes ($\expectationvalue{\mathcal{O}}\sim\absolutevalue{\gamma_{L}}=0$). These zeros can be expressed in a compact form as given in the following equation Maeda:2009vf
@@ 6 Conclusions and discussion
We have developed a holographic model for a type-II s-wave superconductor that incorporates rotating and anisotropic aspects in the black hole background. Using both analytical and numerical approaches, we examined how rotation and anisotropy influence the properties of the superconductor model. Concerning the condensation phenomena, our findings reveal that an increment (or decrement) in the rotation leads to a corresponding reduction (or increment) in the magnitude of the condensation curve, as illustrated in Figure 1. This scaling effect was qualitatively reproduced by the analytical approach, in which the matching method was used, and supported by the numerical approach using the shooting method. Regarding the results for the AC conductivity, we worked out two cases, the isotropic ($z=1$, $\Delta=1$) and the anisotropic one ($z=2$, $\Delta=1$), where novel closed formulas for the conductivity were obtained. The isotropic case remained unaffected by the presence of a nontrivial rotation, indicating that to study the effect induced by this quantity, we must consider the anisotropy of the background. In this regard, the interplay between rotation and anisotropy brought forth a noteworthy behavior not reported in previous holographic models. Concretely, our results have shown that the rotation induces a peak in the low-frequency region of $\text{Re}[\sigma]$, followed by an exponentially vanishing tendency at higher frequencies. Moreover, our findings also revealed that the magnitude of the peak falls monotonically and is shifted to a lower frequency as the rotation increases, see Figure 3(a). Altogether, this behavior aligns with predictions from high-temperature superconductor models and experiments ZIMMERMANN199199 ; PhysRevB.54.700 ; Dressel2013-mq ; Michon2023-lq ; Boyack2023-yu . Within this context, the appearance of the peak and the vanishing behavior at high frequencies are attributed to the effects of quasiparticle damping driven by impurities or defects in the superconducting system. Consequently, our results suggest an interesting connection between the rotation of the anisotropic black hole in our holographic superconductor model and quasiparticle damping in superconducting materials. In relation to the imaginary component $\text{Im}[\sigma]$, the effect of the rotation is manifested in increasing the amplitude of the curve at higher frequencies as the rotation parameter increases, see Figure 3(b).
@@
Furthermore, we considered the superconducting system subject to a uniform magnetic field. We employed the matching method to derive an analytical expression for the upper critical magnetic field, the threshold below which the vortex phenomenon occurs. The results revealed how this upper critical magnetic field gets affected by the dynamical scaling exponent $z$ and the rotation of the background $a$. In both cases, the effect is no more than re-scaling the upper critical magnetic field, as illustrated in Figure 5.
@@ Acknowledgements.
All the authors are grateful to Manuel de la Cruz for enriching discussions. The authors acknowledge financial support from FORDECYT-PRONACESCONACYT Grant No. CF-MG-2558591, CONAHCYT Grant No. A1-S-38041, and VIEP-BUAP Grant. JAHM also acknowledges support from CONAHCYT through a PhD Grant No. 750974. DFHB and JAMZ are also grateful to CONAHCYT for a Estancias posdoctorales por México Grant No. 372516 and project 898686, respectively.

### v2

@@ 5 Turning on a uniform external magnetic field
with $\zeta=-\dfrac{a}{\Xi}\,\rho$. Regular boundary conditions—both asymptotically and at the horizon—constrain the parameters in the following manner: the critical exponent must lie within the interval $1\leq z<3$, while the constants are fixed as $\rho=r_{h}^{3-z}\mu$ and $\zeta=r_{h}^{3-z}\nu$.
@@
$\Biggl{[}$ $u^{z+2}\partial_{u}\Biggl{(}\dfrac{f}{u^{z+2}}\partial_{u}\Biggr{)}+\quantity(\dfrac{\ell u}{r_{h}})^{2(z-1)}\dfrac{(\ell^{2}\Xi A_{t}^{(0)}+aA_{\phi}^{(0)})^{2}}{r_{h}^{2}f}-\ell^{2}\Big{(}\dfrac{m^{2}}{u^{2}}+\dfrac{(aA_{t}^{(0)}+\Xi A_{\phi}^{(0)})^{2}}{r_{h}^{2}}\Big{)}\Biggr{]}\Phi$
@@ 5.1 The external magnetic field
Upon a brief inspection of Eq. (42), it becomes evident that the external magnetic field plays a significant role in the dynamics of the radial mode. As we explore this effect further, the requirement of a consistent solution leads to a nontrivial relation between the upper critical magnetic field and the black hole parameters.
According to the standard holographic prescription, the radial component of the scalar field near the boundary admits the asymptotic expansion
@@
where the scaling dimensions are given by $\Delta_{\pm}=\frac{1}{2}\left(z+3\pm\sqrt{(z+3)^{2}+4\ell^{2}m^{2}}\right)$. Without loss of generality, we impose the boundary condition $J_{+}=0$, identifying $J_{-}=J$ with scaling dimension $\Delta_{-}=\Delta$.
On the other hand, we propose a near-horizon Taylor expansion of $R(u)$ around $u=1$
@@
By evaluating the radial equation (42) near the horizon, the expansion coefficients can be algebraically determined in terms of $R(1)$. The first two coefficients take the form
@@
Once again, the matching method proves essential. Ensuring continuity and regularity requires a smooth matching between (43) and (44), and their derivatives, within a specific region of the bulk domain, say $u=u_{m}$. This matching can be achieved at the expense of the following algebraic constraints
@@
$\times\Biggl{\{}\ell^{2}m^{2}\Big{[}\ell^{2}m^{2}+2(z+3)\Big{]}+\dfrac{\ell^{4}B}{r_{h}^{2}}\Big{[}2\ell^{2}m^{2}+\dfrac{\ell^{4}B}{r_{h}^{2}}\Big{]}-\quantity(\dfrac{\ell}{\Xi})^{2}\quantity(\dfrac{\ell}{r_{h}})^{2z}[{A_{t}^{(0)}}^{\prime}(1)]^{2}\Biggr{\}}$
$\times\,R(1)(u_{m}-1)^{2},$ (46a)
@@
A particularly important implication emerges from evaluating (46) near the upper critical magnetic field, $B\approx B_{c2}$, in accordance with the perturbative assumption $\epsilon\ll 1$. By substituting the expression for $A_{t}^{(0)}{}^{\prime}(1)=(z-3)\rho/r_{h}^{3-z}$, obtained from  (38), we are left with a consistency condition. This condition can be interpreted as an analytical expression for the upper critical magnetic field in terms of the critical temperature, namely,
@@
A quick check-up reveals that $B_{c2}$ goes as predicted by the Ginzburg-Landau theory near the critical temperature, with $B_{c2}\propto(1-T/T_{c})$. This behavior is illustrated in Figure 5, which also shows the effects of rotation and anisotropy on $B_{c2}$, with the scaling dimension set to $\Delta=1$. As shown in the figure, increasing the dynamical exponent $z$ leads to a decrease in the critical magnetic field, suggesting that anisotropic scaling makes the system more susceptible to the suppression of superconductivity. In contrast, the rotation parameter contributes to an increase in the critical magnetic field, implying that a stronger external magnetic field is required to destroy the superconducting phase.
Figure 5: Upper critical magnetic field as function of temperature. The left panel shows the behavior of the upper critical magnetic field for fixed parameters $\Delta=1$, $z=2$, $\ell=1$, and $u_{m}=0.8$ while adopting different values of the rotation parameter $a$. In contrast, the right panel displays the corresponding behavior for $\Delta=1$, $a=1$, $\ell=1$, and $u_{m}=0.8$ considering different values of the critical exponent $z$.
@@ 5.2 The vortex configuration
Let us now transition to a more phenomenological setting, aimed at modeling vortex lattice deformations in experimentally realizable superconductors. In particular, we explore whether the anisotropic holographic setup developed above can reproduce lattice transitions observed in materials such as LiFeAs, following the investigations carried out in previous works Maeda:2009vf ; Guo:2014wca ; Herrera-Mendoza:2022whz .
@@
[…]
@@
Here it is worth noticing that the linear behavior that the $\beta$ parameter presents with respect to the external magnetic field $B$ in the LiFeAs superconductor, contrasts with the quadratic dependence that this parameter displays for the FeSe superconducting material Herrera-Mendoza:2022whz . This difference indicates that the dependence of this parameter on the external magnetic field $\beta(B)$ is determined by the intrinsic properties of diverse superconductors.
@@ 6 Conclusions and discussion
We have developed a holographic model for a type-II s-wave superconductor that incorporates rotating and anisotropic aspects in the black hole background. Using both analytical and numerical approaches, we examined how rotation and anisotropy influence the properties of the superconductor model. Concerning the condensation phenomena, our findings reveal that an increment (or decrement) in the rotation leads to a corresponding reduction (or increment) in the magnitude of the condensation curve, as illustrated in Figure 1. This scaling effect was qualitatively reproduced by the analytical approach, in which the matching method was used, and supported by the numerical approach using the shooting method. Regarding the results for the AC conductivity, we worked out two cases, the isotropic ($z=1$, $\Delta=1$) and the anisotropic case ($z=2$, $\Delta=1$), where novel closed formulas for the conductivity were obtained. The isotropic case remained unaffected by the presence of a nontrivial rotation, indicating that to study the effect induced by this quantity, we must take into account the anisotropic properties of the background. In this regard, the interplay between rotation and anisotropy brought forth a noteworthy behavior not reported in previous holographic models. Concretely, our results have shown that the rotation induces a peak in the low-frequency region of $\text{Re}[\sigma]$, followed by an exponentially vanishing tendency at higher frequencies. Moreover, our findings also revealed that the magnitude of the peak falls monotonically and is shifted to a lower frequency as the rotation increases, as shown in the left panel of Figure 3. Altogether, this behavior aligns with predictions from high-temperature superconductor models and experiments ZIMMERMANN199199 ; PhysRevB.54.700 ; Dressel2013-mq ; Michon2023-lq ; Boyack2023-yu . Within this context, the appearance of the peak and the vanishing behavior at high frequencies are attributed to the effects of quasiparticle damping driven by impurities or defects in the superconducting system. Consequently, our results suggest an interesting connection between the rotation of the anisotropic black hole in our holographic superconductor model and quasiparticle damping in superconducting materials. In relation to the imaginary component $\text{Im}[\sigma]$, the effect of the rotation is manifested in increasing the amplitude of the curve at higher frequencies as the rotation parameter increases, depicted in the right panel of Figure 3.
@@
Furthermore, we considered the superconducting system subject to a uniform magnetic field. We employed the matching method to derive an analytical expression for the upper critical magnetic field, the threshold below which the vortex phenomenon occurs. The results reveal that the upper critical magnetic field is affected by both the dynamical scaling exponent $z$ and the rotation parameter $a$. Specifically, increasing $z$ diminishes the value of the critical magnetic field required to destroy the superconducting state, while increasing $a$ enhances it, see Figure 5. These findings imply that anisotropic scaling tends to weaken superconductivity by making it more sensitive to external magnetic fields, whereas rotation strengthens it in the sense that stronger external magnetic fields are needed to destroy the superconducting state.
@@ Acknowledgements.
All the authors are grateful to Manuel de la Cruz for enriching discussions. The authors acknowledge financial support from FORDECYT-PRONACESCONACYT Grant No. CF-MG-2558591, CONAHCYT Grant No. A1-S-38041, and VIEP-BUAP Grant. JAHM acknowledges support from CONAHCYT through PhD Grant No. 750974 and Postdoctoral support under project No. CF-MG-2558591. DFHB and JAMZ are also grateful to CONAHCYT for a Estancias posdoctorales por México Grant No. 372516 and project 898686, respectively.
Appendix A Analytical study of the condensation operator
Here we want to find an analytic expression for the expectation value of the condensation operator in terms of the black hole parameters. To achieve this, we solve the system (12) using the matching method Zhao:2013pva , which involves determining the solutions at the horizon ($u=1$) and the asymptotic boundary ($u=0$) and matching them at an intermediate point within the bulk $0<u_{m}<1$.
To find the horizon solutions, we propose the following Taylor series expansions of the fields around $u=1$
$\psi(u)$ $=\psi(1)+\psi^{\prime}(1)\quantity(u-1)+\dfrac{\psi^{\prime\prime}(1)}{2}\quantity(u-1)^{2}+\cdots\,$, (60a)
$A_{t}(u)$ $=A_{t}(1)+A_{t}^{\prime}(1)\quantity(u-1)+\dfrac{A_{t}^{\prime\prime}(1)}{2}\quantity(u-1)^{2}+\cdots\,$, (60b)
$A_{\phi}(u)$ $=A_{\phi}(1)+A_{\phi}^{\prime}(1)\quantity(u-1)+\dfrac{A_{\phi}^{\prime\prime}(1)}{2}\quantity(u-1)^{2}+\cdots\,.$ (60c)
In these expressions the regularity of the gauge field at the horizon leads to the boundary conditions
$A_{t}(1)=0,\qquad A_{\phi}(1)=0,$ (61)
while $\psi(1)$ is keep finite. The remaining coefficients are determined by making the expansions (60) consistent with the field equations (12). After some algebraic computations, we find
Note that the higher-order terms are also expressed in terms of the lowest-order ones. On the other hand, the regular solutions at the asymptotic boundary read
$\psi=J\,u^{\Delta},\quad A_{t}=\mu-\rho\quantity(\dfrac{u}{r_{h}})^{3-z},\quad A_{\phi}=\nu-\zeta\quantity(\dfrac{u}{r_{h}})^{3-z},$ (63)
where we identify $J$ as related to the expectation value of the condensation operator, $J=\expectationvalue{\mathcal{O}}/{\sqrt{2}r_{h}^{\Delta}}$, and $\Delta$ as the scaling dimension of the operator $\mathcal{O}$.
Thus far we have found expressions for both regimes, near horizon $(u=1)$ and asymptotic $(u=0)$ solutions. In what follows, we require that there is an intermediate point, $u=u_{m}$, within the bulk, where the asymptotic solutions, (63) and (60), match and their first derivatives as well. This assumption leads us to four independent relations
From (64a) and (64b) we arrive at
where
Similarly, from (64d) we obtain
At this point, let us recall that the black hole temperature is given by Eq. (5) with $d=5$, taking the form
$T=\dfrac{1}{4\pi}\dfrac{(z+3)r_{h}^{z}}{\ell^{z+1}\Xi},$ (68)
which allows us to rewrite Eq. (67) as an explicit function of the black hole temperature
$\psi(1)^{2}=\dfrac{(z+3)}{2(1-u_{m})}\quantity[\dfrac{\ell^{z}(3-z)}{\gamma\,\Xi^{1+3/z}}\quantity(\dfrac{z+3}{4\pi\ell^{z+1}})^{3/z}\dfrac{\rho\,u_{m}^{2-z}}{T^{3/z}}-(z-2)(1-u_{m})-1],$ (69)
which, together with Eq. (65a), provides a concrete expression for the boundary coefficient $J$ in terms of the black hole temperature.